if 3p^2=5p+2 and 3q^2=5q 2, where p=!q then pq=?.............................

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Solve this quadraic first,
3p^2=5p+2
3p^2 – 5p - 2  = 0
3p^2 -6p + p – 2 = 0
3p(p – 2) + (p-2) = 0
p =2  and p = - 1/3
For  p=!q
p = 2 and q = 2
therfore pq = 2*2 = 4